OpenStudy (anonymous):

Blueberry high fiber muffins contain 51% dietary fiber. If a package with a net weight of 12 oz. contains 6 muffins. how many grams of fiber are in each muffin?

OpenStudy (greatlife44):

step one: we need to find out how many ounces are in a single muffin. We know that we have in a total 6 muffins and the total weight is 12 ounces. \[\frac{ 12 ounces }{ 6 muffins } = 2 oz \]

OpenStudy (greatlife44):

Now, we know that each muffin is 2 ounces. that's great. now we must find out what percentage is of this muffin is high fiber. we know that each muffin contains 51% dietary fiber. we know a muffin is 2 ounces We can set up a proportion where x is the weight of fiber. and then we cross multiply \[\frac{ x }{ 2oz } = \frac{ 51 }{ 100 }\] by cross multiplying meaning multiplying x times 100 and 51 times 2 we get this \[100x = 102 \] we divide both sides by 100 \[102/100 = 1.02 = x \] we get x is equal to 1.02 oz. so each muffin contains 1.02 ounces but we were asked how many grams are in each muffin.

OpenStudy (greatlife44):

Doing a simple search we find that \[\frac{ 1 ounce }{ 28.3grams }\] Now to convert ounces to grams we must do something called dimensional analysis; we set it up so that the number we are multiplying by in this case ounces, that ounces is in the denominator in our conversion factor. \[1.02 ounces* (\frac{ 28.3grams }{ 1 ounce } ) = 28.9 grams\]

OpenStudy (greatlife44):

@joyciee91

OpenStudy (anonymous):

converting ounces to grams, and then multiplying by the percent of fiber could be an option too!